# Stick to your Role! Stability of Personal Values Expressed in Large Language Models

### v1

@@ 3.1 Models
The LLaMa-2 (?, ?) family contains models with 7, 13 and 70 billion parameters. It also includes “chat” versions, which were fine-tuned on instructions and with RLHF (?, ?). The Mistral (?, ?) family contains base and instruction fine-tuned models with 7 billion parameters. Zephyr (?, ?) also belongs in this family as a DPO (?, ?) tuned version of the base Mistral model. The Mixtral (?, ?) family contains base and “instruct” models with 46.7 billion parameters. The “instruct” version was trained by supervised fine-tuning and DPO (?, ?). We consider these two models and their 4-bit quantized versions. The Phi (?, ?) family contains smaller base models, of which we consider two models with 1.3 and 2.7 billion parameters. From the Qwen (?, ?) model family we consider base models with 7, 14, and 74 billion parameters.
((a)) Rank-order value stability for LLMs simulating fictional characters.
((b)) Rank-order value stability for LLMs simulating real-world personas.
((c)) Rank-order donation stability on a downstream behavioral task for LLMs simulating fictional characters.
Figure 4: Rank-order stability ($Mean\pm SI(\alpha=0.05)$) of personal values and donations exhibited by LLMs following conversations on different topics. Consistent trends are visible: Mixtral, Qwen, and Mistral model families are the most stable, compared to LLaMa-2 and Phi families. All models exhibit lower than human stability, despite the comparison being skewed in their favor.
Figure 5: Ipsative value stability ($Mean\pm SI(\alpha=0.05)$) of LLMs without the persona setting instructions. Mistral-Instruct-v0.1 and Qwen-72B models show the highest stability. Mixtral, Mistral and Qwen families are more stable. Human change is shown for reference, but no strong conclusions can be made because the comparison is skewed in the LLMs’ favor.
@@ 3.2 How do different models and model families compare in terms of expressed value stability?
We evaluate the Rank-Order stability of LLMs instructed to simulate various personas. Figure 4 compares models’ value stability of two simulated populations: fictional characters (Fig. 3(a)) and real-world personas (Fig. 3(b)). On fictional characters the most stable model is Mixtral-Instruct-v1 ($r=0.42$), which is followed by its 4-bit quantized version ($r=0.32$), Mistral-Instruct-v2 ($r=0.31$), and Qwen-72B ($r=0.26$). A similar trend is observed on real-world personas, however, Qwen-72B ($r=0.5$) matches the stability of Mixtral-Instruct-v1 ($r=0.51$). More generally, we observe consistent trends in terms of model families in both simulated populations: Mixtral, Mistral and Qwen families show more stability than LLaMa-2 and Phi.
Figure 5 compares the Ipsative stability of LLMs without instructing them to simulate any particular persona. While similar trends of models are observed to those in the Rank-Order experiments, the models are less polarized. While Mixtral-Instruct-v0.1 ($r=0.84$), its 4-bit quantized version ($r=0.8$), and Qwen-72B ($r=0.78$) are again the most stable models, zephyr-7b-beta ($r=0.62$) is more stable than Mistral-Instruct-v2 ($r=0.45$). Furthermore, compared to the previous experiment, stability is also observed in LLaMa-2-70b-chat ($r=0.47$) and to a lesser extent in Phi-2 ($r=0.29$), LLaMa-2-70b ($r=0.17$), and Qwen-7B ($r=0.22$). The most stable model families are again Mixtral, Mistral and Qwen.
Instruction or chat fine-tuning seems to be beneficial for Ipsative stability, as every tuned model in Fig. 5 is more stable than its base version. This effect is not as conclusive for Rank-order stability. As fine-tuning adapts the model towards instruction following, dialogues (chat textual format), and answering questions, it is expected to increase stability. However, it also often includes “aligning” the model by making it less prone to exhibit unwanted behavior, which can have a detrimental effect on simulating some personas such as villains. We hypothesize that this is the reason why we observe a consistent effect only on Ipsative stability.
@@
To get a more intuitive impression of the observed stability levels, we extract data from two longitudinal studies on humans. ? (?) followed 20-year-olds for eight years and ? (?) followed 10-year-old for 2 years (these changes are denoted by horizontal lines in figures 4 and 5). It is important to note that this comparison is skewed in the LLMs favor. It is easier for LLMs to show stability in the following ways: 1) human value changes were caused by much more drastic circumstances (years of development compared to topic change in LLMs) 2) the human population was more unstable (10-year-old and 20-year-olds compared to well-established fictional characters or real-world personas). Therefore, an argument can only be made in one direction: if some models show lower stability than that observed in humans, those models can be said to exhibit subhuman value stability.
Figures 3(a) and 3(b) show that all models, when instructed to simulate various personas, exhibit much lower Rank-order stability than that observed in human populations ($r=0.57$ for ages 10 to 12, and $r=0.66$ for ages 20 to 28). The fact that LLMs show lower stability despite the comparison being skewed in their favor shows that LLMs exhibit sub-human value stability and are significantly more susceptible to unexpected context changes. These results motivate reserach on LLMs focused on simulating populations.
Figure 5 shows the Ipsative stability of models that were not instructed to simulate a persona. Both Mixtral-Instruct models ($r=0.84$ and $r=0.8$), Qwen-72B ($r=0.78$), and zephyr-7b-beta ($r=0.62$) do not exhibit lower stability than that observed in humans ($r=0.66$ for ages 10 to 12, and $r=0.59$ for ages 20 to 28)). Crucially, as discussed above, this does not imply that those models show human-level value stability, rather, the only insight is that other models show very low Ipsative stability.
Figure 6: Rank-order value stability ($Mean\pm SI(\alpha=0.05)$) of the Mixtral-Instruct model simulating fictional characters. Stability decreases with longer simulated conversations.
Figure 7: Ipsative value stability ($Mean\pm SI(\alpha=0.05)$) of LLMs with (Mixtral-Instruct) and without persona setting instructions. All models retain the same stability level in longer conversations.
@@ 3.4 Can LLMs keep coherent value profiles over longer conversations?
In the previous experiment, models were let to exchange $n=3$ messages (not including the manually set first and last interlocutor’s messages). Here, we evaluate the effect of simulated conversation length ($n$) on value stability.
@@
Figure 6 shows the effect of simulated conversation length on Rank-order stability expressed by the Mixtral-Instruct model instructed to simulate fictional characters. Due to computational constraints (evaluating a model on one population requires 300k queries), we conduct this experiment only with Mixtral-Instruct (the most stable model from Fig. 3(a)), and only on one population. We can see that, even for this most stable model, stability diminishes as conversations get longer. This experiment highlights the limitations of LLMs in maintaining coherent interpersonal value profiles over longer conversations.
@@
Figure 7 shows the effect of conversation length on Ipsative stability. We compare the most stable models from Fig. 5 without persona instructions, and Mixtral-Instruct with instructions to simulate fictional characters. Ipsative stability remains stable regardless of the simulated conversation length for all models. Mixtral-Instruct with persona instruction ("Mixtral-Instruct (fict. char.)"), while highly stable, is less stable than uninstructed Mixtral-Instruct ("Mixtral-Intruct"). This implies that Mixtral-Instruct is slightly better adapted for use without the persona instructions.
@@
Mixtral-Instruct with persona instructions exhibits a combination of low Rank-Order stability (Fig. 6) and high Ipsative stability (Fig. 7). This implies that the model is able to maintain characters consistently, however those characters might not be adequately impersonated (e.g. simulating different characters with similar value profiles). This hypothesis is confirmed in Appendix B.2. These results suggest that current LLMs are not well suited for use with persona setting instructions, and motivate future research on LLMs focused on simulating specific personas.
@@
3.5 To what extent do conclusions made with PVQ transfer to a downstream behavioral task?

### v4

@@ 4.1 Models
We consider 21 LLMs from the following six model families: LLaMa-2 [41], Mistral [44], Mixtral [45], Phi [46], Qwen [47], and GPT-3.5 [48]. The LLaMa-2 family contains models with 7, 13 and 70 billion parameters (“llama_2_[7|13|70]b”) trained with 2T tokens. It also includes “chat” versions (“llama_2_[7|13|70]b_chat”), which were fine-tuned on instructions and with RLHF [49]. The Mistral family contains base (“Mistral-7B-v0.1”) and instruction fine-tuned models (“Mistral-7B-Instruct-v0.[1|2]”) with 7 billion parameters. Zephyr [50] (“zephyr-7b-beta”) also belongs in this family as a DPO [51] tuned version of the base Mistral model. The Mixtral family contains base (“Mixtral-8x7B-v0.1”) and “instruct” (“Mixtral-8x7B-Instruct-v0.1”) models with 46.7 billion parameters. Those are Mixture-of-Experts models, which means that only 12.9 billion parameters are used per token. The “instruct” version was trained by supervised fine-tuning and DPO. We consider these two models and their 4-bit quantized versions. The Phi family contains smaller base models, of which we consider two models with 1.3 and 2.7 billion parameters (“phi-[1|2]”). From the Qwen family we consider base models with 7, 14, and 74 billion parameters (“Qwen-[7|14 |74]B”), which were trained on 2.2T 2.4T 3T tokens, respectively. From the GPT-3.5 family, we consider the latest two versions: from January 2024 (“gpt-3.5-turbo-0125”) and from October 2023 (“gpt-3.5-turbo-1106”).
@@ 4.2 How do different models and model families compare in terms of expressed value stability?
Rank-order stability
We evaluate the Rank-order stability of LLMs instructed to simulate various personas. Fig 4 compares models’ value stability of two simulated populations: fictional characters (Fig 4(a)) and real-world personas (Fig 4(b)). Statistical analyses for Figs 4(a) and 4(b) are shown in Appendix E in Figs 16 and 17, respectively. On fictional characters (Fig 4(a)) the most stable model is Mixtral-8x7B-Instruct-v0.1 ($r=0.43$), which is followed by its 4-bit quantized version ($r=0.3$), Mistral-7B-Instruct-v0.2 ($r=0.28$), Qwen-72B ($r=0.24$), gpt-3.5-turbo-1106 ($r=0.20$), and gpt-3.5-turbo-0125 ($r=0.15$). A similar trend is observed on real-world personas (Fig 4(a)), however, Qwen-72B ($r=0.46$) approaches the stability of Mixtral-8x7B-Instruct-v0.1 ($r=0.5$). More generally, we observe consistent trends in terms of model families in both simulated populations: Mixtral, Mistral, GPT-3.5, and Qwen families show more stability than LLaMa-2 and Phi.
((a)) Fictional characters simulated population
((b)) Real-world personas simulated population
Figure 4: Rank-order stability with PVQ Rank-order stability ($Mean\pm SE$) of personal values (PVQ) exhibited by simulated participants (fictional characters or real-world personas) following conversations on different topics (correlation of simulated participants’ value expression in different contexts). Consistent trends are visible: Mixtral, Qwen, Mistral, and GPT-3.5 model families are more stable than LLaMa-2 and Phi families. All models exhibit lower than human stability, despite the comparison being skewed in their favor. LLMs are simulating two populations: (a) fictional characters, and (b) real-world personas. For statistical tests, refer to Appendix E (Figs 17 and 16)
Ipsative stability
Fig 5 compares the Ipsative stability of LLMs without instructing them to simulate any particular persona. The statistical analysis is shown in Appendix E in Fig 18. While similar trends of models are observed to those in the Rank-order experiments, the models are less polarized. While Mixtral-8x7B-Instruct-v0.1 ($r=0.84$), its 4-bit quantized version ($r=0.82$), and Qwen-72B ($r=0.73$) are again the most stable models, zephyr-7b-beta ($r=0.62$) is more stable than Mistral-7B-Instruct-v0.2 ($r=0.48$). Furthermore, compared to the previous experiment, stability is also observed in LLaMa-2-70b-chat($r=0.47$) and to a lesser extent in Phi-2 ($r=0.3$), LLaMa-2-70b ($r=0.17$), and Qwen-7B ($r=0.18$). The most stable model families are again Mixtral, Mistral, GPT-3.5 and Qwen.
Instruction or chat fine-tuning seems to be beneficial for Ipsative stability, as every tuned model in Fig 5 is more stable than its base version. This effect is not as conclusive for Rank-order stability. As fine-tuning adapts the model towards instruction following, dialogues (chat textual format), and answering questions, it is expected to increase stability. However, it also often includes “aligning” the model by making it less prone to exhibit unwanted behavior, which can have a detrimental effect on simulating some personas such as villains. We hypothesize that this is the reason why we observe a consistent effect only on Ipsative stability.
Figure 5: Ipsative stability with PVQ Ipsative stability ($Mean\pm SE$) of personal values (PVQ) exhibited by LLMs without the persona setting instructions (correlation of value hierarchies in different contexts). Mistral-7B-Instruct-v0.1 and Qwen-72B models show the highest stability. Mixtral, Mistral, Qwen and GPT-3.5 families are more stable. Human change is shown for reference, but no strong conclusions can be made because the comparison is skewed in the LLMs’ favor. For statistical tests, refer to Appendix E (Fig 18).
@@
To get a more intuitive impression of the observed stability levels, we extract data from two longitudinal studies on humans. [52] followed 20-year-olds for eight years and [53] followed 10-year-olds for 2 years (these changes are denoted by horizontal lines in Fig 4 and Fig 5). It is important to note that this comparison is skewed in the LLMs favor. It is easier for LLMs to show stability in the following ways: 1) human value changes were caused by much more drastic circumstances (years of development compared to topic change in LLMs) 2) the human population was more unstable (10-year-old and 20-year-olds compared to well-established fictional characters or real-world personas). Therefore, an argument can only be made in one direction: if some models show lower stability than that observed in humans, those models can be said to exhibit subhuman value stability.
Fig 4 shows that all models, when instructed to simulate various personas, exhibit much lower Rank-order stability than that observed in human populations ($r=0.57$ for ages 10 to 12, and $r=0.66$ for ages 20 to 28). The fact that LLMs show lower stability despite the comparison being skewed in their favor shows that LLMs exhibit sub-human value stability and are significantly more susceptible to unexpected context changes. These results motivate research on LLMs focused on simulating populations.
Fig 5 shows the Ipsative stability of models that were not instructed to simulate a persona. Both Mixtral-8x7B-Instruct-v0.1 models ($r=0.84$ and $r=0.82$), Qwen-72B ($r=0.73$), and zephyr-7b-beta ($r=0.62$) do not exhibit lower stability than that observed in humans ($r=0.66$ for ages 10 to 12, and $r=0.59$ for ages 20 to 28). Crucially, as discussed above, this does not imply that those models show human-level value stability, rather, the only insight is that other models show very low Ipsative stability.
@@ 4.4 Can LLMs keep coherent value profiles over longer conversations?
In the previous experiment, models were let to exchange $n=3$ messages (not counting the manually set first and last interlocutor’s messages). Here, we evaluate the effect of increased simulated conversation length ($n$) on value stability.
@@
Fig 6 shows the effect of simulated conversation length on Rank-order stability expressed by the Mixtral-8x7B-Instruct-v0.1 model instructed to simulate fictional characters. Due to computational constraints (evaluating a model on one population requires 300k queries), we conduct this experiment only with Mixtral-8x7B-Instruct-v0.1 (the most stable model from Fig 4), and only on one population (fictional characters). We can see that, even for this most stable model, stability diminishes as conversations get longer. It gradually diminishes from $r=0.42$ $(n=3)$ to $r=0.15$ $(n=43)$. We can also see that the stability seems to converge after 35 messages with only a slight drop from 35 to 43 simulated messages ($r=0.166$ to $r=0.162$). This experiment highlights the limitations of LLMs in maintaining coherent interpersonal value profiles over longer conversations.
Figure 6: Rank-order stability with longer conversations Rank-order value stability ($Mean\pm SE$) following conversations of different length for the Mixtral-8x7B-Instruct-v0.1 model simulating fictional characters (correlation of simulated participants’ value expression in different contexts). Stability decreases with longer simulated conversations.
Ipsative stability
@@
Fig 7 shows the effect of conversation length on Ipsative stability. We compare the most stable models from Fig 5 without persona instructions, and Mixtral-8x7B-Instruct-v0.1 with instructions to simulate fictional characters. Ipsative stability remains stable regardless of the simulated conversation length for all models. Mixtral-8x7B-Instruct-v0.1 with persona instructions (“Mixtral-8x7B-Instruct-v0.1 (fict. char.)”), while highly stable, is less stable than the uninstructed model (“Mixtral-8x7B-Instruct-v0.1”). This implies that Mixtral-8x7B-Instruct-v0.1 is slightly better adapted for use without the persona instructions.
Figure 7: Ipsative stability with longer conversations Ipsative value stability ($Mean\pm SE$) of LLMs with (fict. char.) and without persona setting instructions (correlation of value hierarchies in different contexts). All models retain the same stability level in longer conversations.
@@
Mixtral-8x7B-Instruct-v0.1 with persona instructions exhibits a combination of decreasing low Rank-order stability (Fig 6) and high Ipsative stability (Fig 7). This implies that as conversations get longer, the simulated persona drifts towards a similar value profile. This hypothesis is confirmed in Appendix D.2. These results suggest that current LLMs are not well suited for use with persona setting instructions, and motivate future research on LLMs focused on simulating specific personas. We hypothesize this to be a consequence of instruction fine-tuning, which is currently biased towards assistant-like chatbots.
@@
4.5 To what extent do conclusions made with PVQ transfer to downstream behavioral tasks?
((a)) Religion task
((b)) Donation task
((c)) Stealing task
Figure 8: Rank-order stability on downstream tasks Rank-order stability ($Mean\pm SE$) on behavioral downstream tasks of various LLMs (correlation of simulated participants’ behavior in different contexts). Three downstream tasks are shown: (a) Religion. (b) Donation, and (c) Stealing. For statistical test, refer to Appendix E (Figs 19, 20, and 21). Consistent trends with the PVQ experiments (Fig 4) are visible. Mixtral, Qwen, Mistral, and GPT-3.5 model families are the most stable, compared to LLaMa-2 and Phi families. Mixtral-8x7B-Instruct-v0.1, Mistral-7B-Instruct-v0.2, and Qwen-72B are the most stable models. Trends are the most present on the, easiest, Religion task (a) and almost disappear on the, hardest, Stealing task (c).
In this experiment, we study if the conclusions made with the PVQ questionnaire transfer to a downstream behavioral task, i.e. if models that exhibited more stable value profiles also exhibit more stable behavior on a downstream task. We construct three downstream tasks: Donation, Religion, and Stealing. Here we briefly describe them and give more details in appendix C.